# Advances in Embodied Navigation Using Large Language Models: A Survey

### v4

@@ Abstract
In recent years, the rapid advancement of Large Language Models (LLMs) such as the Generative Pre-trained Transformer (GPT) has attracted increasing attention due to their potential in a variety of practical applications. The application of LLMs with Embodied Intelligence has emerged as a significant area of focus. Among the myriad applications of LLMs, navigation tasks are particularly noteworthy because they demand a deep understanding of the environment and quick, accurate decision-making. LLMs can augment embodied intelligence systems with sophisticated environmental perception and decision-making support, leveraging their robust language and image-processing capabilities. This article offers an exhaustive summary of the symbiosis between LLMs and embodied intelligence with a focus on navigation. It reviews state-of-the-art models, research methodologies, and assesses the advantages and disadvantages of existing embodied navigation models and datasets. Finally, the article elucidates the role of LLMs in embodied intelligence, based on current research, and forecasts future directions in the field. A comprehensive list of studies in this survey is available at https://github.com/Rongtao-Xu/Awesome-LLM-EN.
keywords:
@@
journal: Engineering Applications of Artificial Intelligence
1 Introduction
@@
To address these challenges, researchers have employed a diverse array of methods such as machine learning [1], deep learning [2] and evolutionary algorithms [3] [4], as well as the addition of other vision-language models to enhance the multimodal data fusion capabilities of LLMs. These methodologies aim to develop navigation agents capable of learning from experience and continually improving their performance. In this paper, we review existing models that have employed LLM-based agents for navigation tasks. These studies have demonstrated that LLM-based agents perform excellently in path planning, environmental understanding, and dynamic adjustment tasks. The success of these studies underscores the utility of LLM-based agents as valuable tools for navigation.
@@
    Dataset Analysis: The article thoroughly analyzes commonly used datasets in the navigation field, discussing their applicability, issues, and limitations. This detailed analysis guides researchers in selecting appropriate datasets and offers practical advice for their application, enhancing the article’s practical relevance.
@@
2 Background
2.1 Large Language Models
The development of LLMs represents a significant milestone in the domains of Natural Language Processing (NLP) [8] and Machine Learning. LLMs are machine learning models specifically designed to process and generate natural language [9]. They are typically trained on extensive text data to capture the grammatical structures, inter-word relationships, and contextual nuances of language. These models predominantly utilize the Transformer architecture, which is distinguished by its self-attention mechanism that effectively addresses long-range dependencies. By employing pre-training on broad corpora to learn fundamental language patterns followed by fine-tuning on specialized datasets for specific tasks, LLMs are adapted for diverse applications. Additionally, these models are capable of not only understanding complex textual information but also generating coherent and fluent text outputs, demonstrating their high efficacy in natural language generation, translation, summarization, and question answering. Key technical features of LLMs include:
  1. 1.
    Word Embedding: Models process text by converting words into numerical vectors, known as word embeddings, which encapsulate the semantic and syntactic properties of vocabulary, laying the groundwork for further language processing.
  2. 2.
    Layers and Heads: The Transformer architecture comprises multiple layers, each containing several ’heads’ (Attention Heads) that process different types of information in parallel, enhancing the model’s capacity to handle information.
  3. 3.
    Sequence Processing: Unlike Recurrent Neural Networks, Transformers can process entire input sequences in parallel, significantly improving efficiency and speed.
Tracing the evolution of these models requires examining the early stages of NLP and machine learning, where methods such as Bag-of-Words (BoW) [10] were prevalent for text representation but limited in capturing word order and contextual nuances.
Before the mainstream adoption of deep learning, simpler algorithms like N-grams and decision trees held sway in a variety of language-processing tasks. N-grams served as foundational elements for text and language data analysis by breaking down text into sequences of N contiguous words or characters. Although efficient and straightforward, N-grams suffer from limitations such as failing to capture long-distance dependencies and requiring a large parameter space for substantial N-values.
Decision trees have been employed for classification and regression tasks. These trees recursively partition the dataset, where each internal node denotes a feature, each branch symbolizes a decision rule, and each leaf node represents the final output. While decision trees offer ease of interpretation and handle irrelevant features effectively, they are susceptible to overfitting and struggle with non-linear, high-dimensional data.
The advent of word embedding models around 2013, notably Word2Vec and GloVe, signaled a transformative shift in NLP. These models allowed for words to be mapped into a continuous vector space, thereby capturing intricate word relationships with remarkable effectiveness. With the advent of word embedding techniques , the focus shifted toward incorporating these pre-trained word vectors into sequence models like Recurrent Neural Networks (RNNs) and LSTMs. RNNs are neural architectures specifically designed to handle sequence data by utilizing the temporal context from prior elements in a sequence for processing subsequent inputs. Nevertheless, during their training, RNNs are plagued by the vanishing and exploding gradient problems, which hinder their capacity to capture long-term dependencies and contribute to training instability.Both models have their merits and drawbacks, but they are now primarily used as baseline models or for feature engineering in the wake of more complex models enabled by deep learning.
To mitigate these drawbacks, specialized RNN variants such as Gated Recurrent Units (GRUs) [11] have gained prominence in sequence data analysis, particularly within NLP. LSTMs, equipped with gating mechanisms, effectively circumvent the vanishing gradient issue but at the cost of added computational complexity and the potential for exploding gradients. The increased number of parameters also increases the risk of overfitting and extends training time.
Building on word embeddings and sequence models, a new paradigm emerged with the development of end-to-end learning architectures, most notably the Transformer model. Introduced by Google in 2017, the Transformer architecture eschews recurrent and convolutional layers, opting instead for a Self-Attention mechanism to capture sequence dependencies [12]. This design enables parallel data processing, resulting in a significant speed-up over RNNs and LSTMs. The Transformer’s scalability also benefits from its layered architecture, facilitating the addition of layers or parameters to improve performance.
Following the Transformer, Google’s Vision Transformer (ViT) [13] in 2020 and OpenAI’s CLIP [14] in 2021 marked the model’s extension from NLP to visual processing tasks. Both architectures employ Transformer-based text encoders and demonstrate the model’s versatility across domains.Among the various Transformer-based architectures, Bidirectional Encoder Representations from Transformers (BERT) [15], introduced by Google in 2018, garnered attention for its pre-training approach across diverse NLP tasks. Employing bidirectional encoders, BERT captures both preceding and succeeding contextual information for each word. While effective, BERT’s complexity results in high computational requirements. OpenAI’s introduction of the GPT series [16] [17] [18] has been another milestone, particularly for text generation tasks. These models also exhibit ”zero-shot learning” capabilities, enabling them to perform certain tasks without task-specific training. However, for tasks requiring higher accuracy, fine-tuning is generally necessary.
GPT models have been deployed in a wide array of applications in NLP and artificial intelligence, such as chatbots, automated writing, and question-answering systems, and have opened new avenues in multi-modal learning and code generation.
In summary, LLMs are increasingly shaping various domains, and their ongoing development continues to reveal new potential. LLMs have been widely applied in multiple areas including, but not limited to, natural language processing tasks [9], autonomous driving [5], robotics task planning [19], and knowledge-intensive tasks [20]. Particularly in the fields of autonomous driving and robotics task planning, LLMs help develop precise and efficient action plans by understanding natural language instructions and environmental information [21]. For instance, in the context of self-driving vehicles or mobile robots, LLMs can assist the system in understanding descriptions of destinations and generating route guides to reach specified locations. Moreover, these models can handle user queries, such as information requests about nearby restaurants or specific places, thereby enhancing the user experience of interactive navigation systems. In embodied systems, by integrating LLMs, devices can not only interact more intelligently with users but also adapt and make efficient decisions in a broader range of environments, showcasing the integration and innovation of technology.The widespread deployment and diverse applications of these models not only underscore their immediate utility but also open exciting possibilities for future research.
2.2 Embodied Intelligence
@@
Embodied Intelligence is an emerging field focused on understanding and developing intelligent agents that closely interact with their environment. This field posits that intelligence fundamentally emerges from an agent’s interaction with its surroundings, rather than being an inherent characteristic of isolated, abstract computations. Drawing on various disciplines such as neuroscience, psychology, robotics, and artificial intelligence, Embodied Intelligence seeks to create innovative models and algorithms for simulating intelligent behavior [22].
In the recent developments within the field of Embodied Intelligence, the integration of LLMs with intelligent agents has become a hot topic. This approach leverages the extensive knowledge and excellent NLP capabilities of large language models, translating human instructions into formats interpretable by physically embodied agents. Although large language models excel in natural language processing, they typically lack direct interaction with the physical world, as their training data is predominantly textual. To enhance the efficiency and generalization ability of these agents, many researchers are no longer limiting themselves to using only text as input. Instead, they are integrating images, point clouds [23] [24], and even voice prompts into their projects. Including using finetuned large language models and training with methods such as reinforcement learning [25] [26], these ways undoubtedly align with the concept of Embodied Intelligence and marks a progression towards the integration of multimodal algorithms [27]. Multimodality refers to the integration and utilization of multiple sensory modalities in perception, communication, and cognition. By leveraging multimodal information, such as visual, auditory, and tactile cues, embodied agents can gather a more comprehensive understanding of their environment . This multimodal perception enables them to make sense of complex spatial information, recognize objects and events, and navigate effectively in real-world scenarios [28].
LLMs such as GPT and other multimodal models have demonstrated significant capabilities in processing both images and language, integrating natural language processing and computer vision technologies. These models are trained on vast datasets of images and corresponding textual descriptions, learning to associate visual content with natural language, which allows them to excel in various tasks like image classification and caption generation. Particularly in embodied intelligence systems, such as smart homes, robots, or autonomous vehicles, the image and language processing abilities of LLMs can significantly enhance user interaction. The systems analyze images and leverage language models to understand user needs, enabling smart responses. Some multimodal LLMs process images by feeding them into various image encoders, such as the Vision Transformer [13], and then combine them with relevant context into a sequence of tokens, which are then input into the corresponding LLM for image recognition and processing. Additionally, models like DALL-E allow users to generate new images or edit existing ones by inputting descriptive text [29]. For example, a user might request ”a cat in a spacesuit on the moon,” and the model can generate an image based on this description. LLMs also provide decision support in scenarios requiring real-time information processing and adapt to specific environments or needs through interaction with users. This combination of advanced image processing technology and sophisticated language understanding enables LLMs to be extensively applied in enhancing embodied intelligence systems, offering unprecedented levels of interaction and automation.
Besides, In the realm of Embodied Intelligence, the concept of an intelligent agent is a pivotal element. Control over these agents is divided into high-level and low-level aspects. High-level controls involve task scheduling and strategy development, incorporating techniques like reinforcement learning [30], deep learning [31] [32] [33] [34] [35], and emerging methodologies based on expansive language models. In contrast, low-level control relates to direct commands over the agent’s operational functions, such as position, speed, and force. An agent’s capabilities depend on its design and low-level control parameters [36], while high-level controls orchestrate the execution and approach to task completion [37].
These methodologies have led to a range of distinct applications, including terrain recognition [38], predicting machinery lifespan [39], and emulating gaze mechanisms [40]. These developments not only reflect the dynamic nature of the field but also signify the tangible impact of these controlled embodied agents across various domains. The advancement of intelligent agents hinges on the harmonious integration of high-level and low-level controls. Utilizing superior algorithms and control techniques is essential for developing new agents characterized by greater robustness and broader generalization capabilities. In addition to research on single-agent systems, studies on multi-agent systems focus independently on multi-agent cooperation issues [41] [42] [43]. Their research efforts are instrumental in expanding the range of tasks achievable by agents, enhancing work efficiency, and improving the real-world applicability and generalization of Embodied Agents [44].
To encapsulate, the field of Embodied Intelligence is navigating a transformative journey: starting from reinforcement learning, shifting towards the utilization of LLMs, and currently progressing into the realm of multimodal integration [45]. As a subset of Artificial Intelligence, the advancement of Embodied Intelligence is intrinsically tied to the practical deployment of diverse algorithms within the physical world.
3 LLMs in Embodied Agents
3.1 LLMs for Grounded Language Understanding
@@
Grounded Language Understanding aims to reconcile the abstract symbols processed by language models with concrete entities, actions, or states in the physical or simulated world. This is pivotal for applications requiring real-world interaction, such as robotic control, natural language interfaces, or advanced research in Embodied Intelligence.
@@
Although LLMs have advanced significantly in NLP, they primarily excel in text generation or classification and seldom engage with tangible entities or real-world situations. Their limitations include a lack of foundational knowledge, impairing their ability to process interrelated concepts and function in interactive settings (Mahowald et al. [46]). To mitigate these issues, researchers have employed various strategies, such as 1) fine-tuning pre-trained models, 2) implementing reinforcement learning algorithms for decision-making in complex environments, and 3) devising specialized architectures for multimodal learning.
@@
3.2 LLMs for Few-Shot Planning
@@
LLMs have achieved notable success in NLP and are increasingly being applied to Few-Shot Planning, a machine learning technique designed for effective planning and decision-making in new tasks with minimal sample data. LLMs excel in zero-shot and few-shot learning, allowing for efficient generalization across tasks without the need for task-specific training or limited examples. Employing LLMs for Few-Shot Planning often involves processing natural language queries, enabling the models to generate actionable plans or steps, thus proving invaluable in zero-shot navigation tasks.
@@
As discussed in research by Zhou et al. in ESC [47] and Song et al. in LLM-Planner [48], the effectiveness of navigation tasks is closely related to semantic scene understanding, which relies heavily on Few-Shot Planning. Planners incorporating LLMs are increasingly vital for enabling embodied agents to execute complex tasks in visually rich environments based on natural language directives. The performance of such planners has a direct bearing on the task’s overall success. The role of LLMs as planners in embodied tasks is expanding, thanks in part to the emergence of methodologies like LLM-Planner [48], CoT (Chain-of-Thought)[49], and LLM-DP[50]. These methods complement the already respectable performance of traditional symbolic planners like the Fast-Forward planner [51] and BFS(f) planner [52] in current research.
3.3 Embodied Navigation
@@
In the field of embodied navigation, state-of-the-art models enhance their performance and practicality through various methods. Firstly, these models achieve multimodal integration [45], combining visual, linguistic, and auditory data, using LLMs to process complex language instructions in conjunction with visual systems to accurately interpret and respond to the environment [53]. Secondly, they enhance understanding of environmental context and decision-making through deep learning algorithms [2], leveraging the advanced language comprehension capabilities of LLMs to better grasp task requirements and navigate environmental obstacles [54]. Additionally, models with long-term memory mechanisms can remember past navigation experiences and environmental features, aiding effective path planning in complex or changing environments [54] [55]. These models also possess the ability to interact effectively with users, processing natural language inputs to communicate in a more natural and human-like manner. Lastly, the adoption of reinforcement learning methods allows models to self-optimize and learn through continuous interaction with the environment, continually adjusting behaviors to adapt to new tasks or environmental changes. By integrating these technologies, state-of-the-art embodied navigation models not only improve navigation efficiency but also enhance their ability to cope with complex environments, achieving highly intelligent and adaptable navigation capabilities [6].
3.3.1 LLM-based Model
@@
In recent years, a rapidly expanding array of works has emerged in the field of navigation, each characterized by its unique emphasis and methodologies for performance assessment. Generally, these works employ approaches that fall into two categories, as illustrated in Fig.2 and Fig.3. The first category uses LLMs as planners, directly generating actions and leveraging exploration policies to guide agents. The second category employs LLMs to scrutinize incoming visual or textual data to isolate goal-relevant information, based on which exploration policies subsequently produce suitable actions for agent navigation. An architectural diagram delineating these approaches is presented below. This section endeavors to examine the nuanced features and distinctions among these works, furnishing a comprehensive temporal map of the works under discussion in this chapter, as depicted in Fig.1. In doing so, we aim to offer an exhaustive understanding that will prove invaluable for both scholars and professionals in this field.
Figure 1: This presentation exhibit a temporal map depicting the works of embodied navigation from 2022 to 2023, and we selected 5 typical works to showcase their corresponding framework diagrams. The map illustrates the evolution of major works, offering valuable insights into the advancement of Embodied Agents.
Table 1: Classification of the works described in this section.
@@
Planning CLIP-Nav NavGPT VELMA $A^{2}$av MiC SayNav
Semantic Understanding LM-Nav L3MVN ESC SQA3D
@@
1)LM-Nav [56] is a navigation architecture designed for robots that leverages pre-existing models specialized in language, vision, and action to enable sophisticated interactions with robots via natural language commands. Remarkably, the system eliminates the need for costly supervision and fine-tuning, relying solely on pre-trained models for navigation, image-langu-age correlation, and language modeling. The architecture of LM-Nav consists of three integrated, pre-trained models to ensure precise instruction execution in complex, real-world scenarios. Specifically, the system employs GPT-3 as the LLM, tasked with decoding verbal instructions into a series of textual landmarks. Concurrently, CLIP serves as the VLM, anchoring these textual landmarks to a topological map. Lastly, the Vision-Action Model (VAM) is a self-supervised robotic control model, responsible for utilizing visual data and executing physical actions based on plans synthesized by the LLM and VLM.
@@
Implemented on a real-world mobile robot, LM-Nav has been shown to accomplish long-horizon navigation in intricate, outdoor settings solely based on natural language directives. A salient feature is the lack of model fine-tuning; all three component models are trained on expansive datasets with self-supervis-ed learning objectives and are deployed as-is. The system has demonstrated its ability to interpret and execute natural language instructions across significant distances in complex, suburban terrain while also offering disambiguation in path selection through detailed commands. Such performance metrics underscore LM-Nav’s robust generalization capabilities and its proficiency in navigating complicated outdoor environments.
@@
2)CLIP-NAV [57] introduces an innovative ”zero-shot” navigation framework aimed at solving coarse-grained instruction-following tasks. The architecture is structured to dissect the guidance language into critical keyphrases, visually anchor them, and leverage the resulting grounding scores to direct the CLIP-Nav sequence-to-sequence model in predicting the agent’s subsequent actions. Initially, a keyphrase extraction module isolates salient terms from the given instructions. Following this, a visual grounding module anchors these keyphrases within the environment, thereby generating a set of grounding scores. These scores serve as the basis for the sequence-to-sequence model in CLIP-Nav, which computes the next set of actions based on the agent’s current state and grounding scores.
@@
To augment the efficacy of CLIP-Nav, the model incorporates a backtracking mechanism. This allows the agent to retrace its steps, facilitating revisions to prior decisions. Such a feature is particularly beneficial in the context of coarse-grained instruction-following tasks, wherein corrective backtracking may be essential.
Evaluation metrics confirm the robustness of CLIP-Nav, establishing a zero-shot baseline on the REVERIE task. Remarkably, CLIP-Nav exceeds the performance of the unseen supervised baseline, even without dataset-specific fine-tuning, in both success rate and success weighted by path length. Additionally, the paper introduces a new performance metric, termed Relative Change in Success (RCS), to quantitatively assess generalizability in vision and language navigation tasks. The RCS metric substantiates the superior performance of CLIP-based methodologies over conventional supervised approaches.
Figure 2: The first type employs LLMs as planners that directly generate actions, thereby leveraging exploration policies to control agents. LLMs, by acquiring information from text and other visual models processed through images, perform planning (using a dialogue format as an example here) and hand the actions over to exploration algorithms and agents for navigation to complete tasks.
@@
3)SQA3D [58] (Situated Question Answering in 3D Scenes) introduces a task formulated to assess the scene comprehension aptitude of embodied agents. The task necessitates that the agent garner an exhaustive understanding of its orientation within a 3D environment, guided by a text-based description, and subsequently generate precise answers to questions pertaining to that understanding.

### v5

@@ Abstract
Large Language Models like GPT have gained significant attention for their potential in various practical applications. One notable area of focus is their integration with Embodied Intelligence, particularly in navigation tasks, which require deep environmental understanding and rapid, accurate decision-making. LLMs enhance embodied intelligence systems by providing advanced environmental perception and decision-making support, leveraging their language and image-processing capabilities. This article reviews the symbiosis between LLMs and embodied intelligence, focusing on navigation, and evaluates current models, research methodologies, and datasets. Additionally, it discusses the role of LLMs in the field and forecasts future research directions. A comprehensive list of studies is available at https://github.com/Rongtao-Xu/Awesome-LLM-EN.
Index Terms:
@@
I Introduction
@@
To address these challenges, researchers have employed a diverse array of methods such as machine learning [103], deep learning [148][130] and evolutionary algorithms [52] [118], as well as the addition of other vision-language models to enhance the multimodal data fusion capabilities of LLMs[44]. These methodologies aim to develop navigation agents capable of learning from experience and continually improving their performance[172]. In this paper, we review existing models that have employed LLM-based agents for navigation tasks. These studies have demonstrated that LLM-based agents perform excellently in path planning, environmental understanding, and dynamic adjustment tasks. The success of these studies underscores the utility of LLM-based agents as valuable tools for navigation.
@@
    Dataset Analysis: The article provides a thorough analysis of commonly used datasets in the navigation field, discussing their applicability, issues, and limitations. This detailed analysis helps researchers select appropriate datasets and offers practical advice on their application, thereby enhancing the practical relevance of the article.
@@
II Background
II-A Large Language Models
Figure 1: This presentation exhibit a temporal map depicting the works of embodied navigation from 2022 to 2024, and we selected 5 typical works to showcase their corresponding framework diagrams. The map illustrates the evolution of major works, offering valuable insights into the advancement of Embodied Agents.
The advent of LLMs marks a pivotal achievement in the domains of Natural Language Processing (NLP) and Machine Learning. LLMs are sophisticated machine learning models specifically engineered to process and generate natural language, typically trained on vast text corpora to capture syntactic structures, inter-word relationships, and contextual subtleties [88][72]. These models predominantly utilize the Transformer architecture, which incorporates a self-attention mechanism to adeptly manage long-range dependencies, making them suitable for a wide array of tasks, such as natural language generation, translation, summarization, and question answering.
The Transformer architecture, introduced by Google in 2017 [124], eschews recurrent and convolutional layers in favor of a self-attention mechanism to capture sequence dependencies, thereby dramatically enhancing processing speed over traditional RNNs and LSTMs. Following this, models such as ViT [37] and CLIP [97] expanded the application of Transformer-based models from NLP to visual processing tasks. BERT [35], introduced by Google in 2018, leverages bidirectional encoders to capture both antecedent and subsequent contextual information, while the GPT series, known for its focus on text generation, demonstrates ”zero-shot learning” capabilities, enabling models to execute tasks without task-specific training [98][99][13].
LLMs have found widespread application across various fields, including NLP, autonomous driving, and robotics task planning. In particular, within autonomous driving and robotics task planning, LLMs facilitate the understanding of natural language commands and the generation of precise action plans, thereby enhancing both user experience and operational efficiency [31][162][129]. However, despite their exceptional adaptability and capacity for real-time data processing, LLMs demand substantial computational resources and rely heavily on high-quality, extensive datasets [107].
In conclusion, LLMs exhibit immense potential across a broad range of domains, yet they are confronted with challenges related to data diversity, granular navigation, spatial reasoning, and interaction adaptability. Future advancements are poised to focus on multimodal integration and optimization strategies [160][127].
II-B Embodied Intelligence
@@
Embodied Intelligence is an emerging field that explores intelligent agents interacting with their environment, with the premise that intelligence arises from this interaction rather than abstract computation. This field integrates neuroscience, psychology, robotics, and AI to develop models and algorithms for simulating intelligent behavior [19].
Embodied navigation refers to autonomous movement through perception, decision-making, and action within a physical environment[104]. It requires intelligent agents to understand their surroundings through sensory systems and adapt navigation behaviors based on physical attributes such as size, weight, and mobility [171]. Challenges include dynamic obstacles, terrain variations, limited sensory data, and the need for real-time decision-making [178].
Recent advancements include integrating LLMs with intelligent agents. These models leverage the NLP capabilities of LLMs to translate human instructions into formats interpretable by agents. Researchers are moving beyond text inputs to incorporate images, point clouds [132][131], and voice prompts, aligning with the concept of multimodal algorithms[70]. Multimodal integration allows agents to gather comprehensive environmental data through visual, auditory, and tactile cues, improving navigation in real-world scenarios [144] [113].
Models like DALL-E and Vision Transformers [37] enhance multimodal interaction by processing images and combining them with text for better user interaction. These systems provide decision support, enabling robots to understand user needs and enhance the navigation experience.
In embodied navigation, agents’ control is divided into high-level and low-level aspects. High-level controls focus on task scheduling and strategy development through techniques like reinforcement learning [46], while low-level controls manage direct operations such as speed and position [10]. These methodologies enable applications in terrain recognition [125], machinery lifespan prediction [137], and gaze mechanisms [78], with significant advancements in agent robustness and generalization.
In conclusion, the field of embodied navigation is evolving from reinforcement learning to LLM integration and multimodal approaches [147]. The integration of LLMs and other algorithms continues to enhance intelligent agent capabilities, with a focus on improving adaptability and performance in real-world applications.
Figure 2: The first type utilizes LLMs to analyze incoming visual or textual data to extract goal-relevant information, upon which exploration policies subsequently generate appropriate actions to guide agent movement. LLMs, by acquiring information from text and other visual models processed through images, perform semantic understanding rather than planning. They extract key information sequences such as targets and locations, and hand them over to exploration algorithms. The exploration algorithms generate actions to guide agents in navigation to complete tasks.
III LLMs in Embodied Navigation
Large language models often play two roles in navigation: either as planning models or as information-providing models, as described in this section under ”LLMs for Grounded Language Understanding” and ”LLMs for Few-Shot Planning.”
III-A LLMs for Grounded Language Understanding
@@
Grounded Language Understanding aims to reconcile the abstract symbols processed by language models with concrete entities, actions, or states in the physical or simulated world. This is pivotal for applications requiring real-world interaction, such as robotic control, natural language interfaces, or advanced research in Embodied Intelligence [146].
@@
Although LLMs have advanced significantly in NLP, they primarily excel in text generation or classification and seldom engage with tangible entities or real-world situations. Their limitations include a lack of foundational knowledge, impairing their ability to process interrelated concepts and function in interactive settings (Mahowald et al. [82][17]). To mitigate these issues, researchers have employed various strategies, such as 1) fine-tuning pre-trained models, 2) implementing reinforcement learning algorithms for decision-making in complex environments, and 3) devising specialized architectures for multimodal learning.
@@
Figure 3: The second type employs LLMs as planners that directly generate actions, thereby leveraging exploration policies to control agents. LLMs, by acquiring information from text and other visual models processed through images, perform planning (using a dialogue format as an example here) and hand the actions over to exploration algorithms and agents for navigation to complete tasks.
III-B LLMs for Few-Shot Planning
@@
LLMs excel in zero-shot and few-shot learning, allowing for efficient generalization across tasks without the need for task-specific training or limited examples. Employing LLMs for few-shot planning often involves processing natural language queries, enabling the models to generate actionable plans or steps, thus proving invaluable in zero-shot navigation tasks [74].
@@
As discussed in research by Zhou et al. in ESC [180] and Song et al. in LLM-Planner [116][90], the effectiveness of navigation tasks is closely related to semantic scene understanding, which relies heavily on Few-Shot Planning. Planners incorporating LLMs are increasingly vital for enabling embodied agents to execute complex tasks in visually rich environments based on natural language directives. The performance of such planners has a direct bearing on the task’s overall success. The role of LLMs as planners in embodied tasks is expanding, thanks in part to the emergence of methodologies like LLM-Planner [116], CoT (Chain-of-Thought)[134], and LLM-DP[32]. These methods complement the already respectable performance of traditional symbolic planners like the Fast-Forward planner [50] and BFS(f) planner [75] in current research.
III-C Embodied Navigation
@@
In the field of embodied navigation, state-of-the-art models enhance their performance and practicality through various methods. Firstly, these models achieve multimodal integration [147][158], combining visual, linguistic, and auditory data, using LLMs to process complex language instructions in conjunction with visual systems to accurately interpret and respond to the environment [101][61]. Secondly, they enhance understanding of environmental context and decision-making through deep learning algorithms [148][106], leveraging the advanced language comprehension capabilities of LLMs to better grasp task requirements and navigate environmental obstacles [179][87]. Additionally, models with long-term memory mechanisms can remember past navigation experiences and environmental features, aiding effective path planning in complex or changing environments [179] [96][84]. These models also possess the ability to interact effectively with users, processing natural language inputs to communicate in a more natural and human-like manner[172]. Lastly, the adoption of reinforcement learning methods allows models to self-optimize and learn through continuous interaction with the environment, continually adjusting behaviors to adapt to new tasks or environmental changes [128]. By integrating these technologies, state-of-the-art embodied navigation models not only improve navigation efficiency but also enhance their ability to cope with complex environments, achieving highly intelligent and adaptable navigation capabilities [167][1].
Embodied navigation is crucial for enabling robots to interact effectively with their environments, involving complex tasks like sensory perception, decision-making, and motion planning. Recent research advances have focused on improving path planning algorithms for dynamic and complex environments [117]. [71] introduces adaptive dynamic programming for real-time path adjustments in mobile robots, while [92] applies a memory-based gravity search algorithm for humanoid robot navigation. [81] reviews trajectory planning methods for multi-robot systems, and [53] develops hybrid optimization techniques for autonomous car navigation.
III-C1 LLM-based Model
@@
Generally, LLM-based navigation models employ approaches that fall into two categories, as illustrated in Fig.3 and Fig.2. The first category uses LLMs as planners, directly generating actions and leveraging exploration policies to guide agents. The second category employs LLMs to scrutinize incoming visual or textual data to isolate goal-relevant information, based on which exploration policies subsequently produce suitable actions for agent navigation. An architectural diagram delineating these approaches is presented below. This section endeavors to examine the nuanced features and distinctions among these works, furnishing a comprehensive temporal map of the works under discussion in this chapter, as depicted in Fig.1.
TABLE I: Classification of the works described in this section.
@@
Planning CLIP-Nav NavGPT VELMA MiC SayNav
Semantic Understanding LM-Nav BEVBert ESC SQA3D
@@
a) LM-Nav [109] is a navigation architecture designed for robots that leverages pre-existing models specialized in language, vision, and action to enable sophisticated interactions with robots via natural language commands. Remarkably, the system eliminates the need for costly supervision and fine-tuning, relying solely on pre-trained models for navigation, image-langu-age correlation, and language modeling. The architecture of LM-Nav consists of three integrated, pre-trained models to ensure precise instruction execution in complex, real-world scenarios. Specifically, the system employs GPT-3 as the LLM, tasked with decoding verbal instructions into a series of textual landmarks. Concurrently, CLIP serves as the VLM, anchoring these textual landmarks to a topological map. Lastly, the Vision-Action Model (VAM) is a self-supervised robotic control model, responsible for utilizing visual data and executing physical actions based on plans synthesized by the LLM and VLM.
@@
Implemented on a real-world mobile robot, LM-Nav has been shown to accomplish long-horizon navigation in intricate, outdoor settings solely based on natural language directives. A salient feature is the lack of model fine-tuning; all three component models are trained on expansive datasets with self-supervised learning objectives and are deployed as-is. The system has demonstrated its ability to interpret and execute natural language instructions across significant distances in complex suburban terrain, while also offering disambiguation in path selection through detailed commands. Such performance metrics underscore LM-Nav’s robust generalization capabilities and its proficiency in navigating complicated outdoor environments.
@@
b) CLIP-NAV [36] introduces an innovative ”zero-shot” navigation framework aimed at solving coarse-grained instruction-following tasks. The architecture is structured to dissect the guidance language into critical keyphrases, visually anchor them, and leverage the resulting grounding scores to direct the CLIP-Nav sequence-to-sequence model in predicting the agent’s subsequent actions. Initially, a keyphrase extraction module isolates salient terms from the given instructions. Following this, a visual grounding module anchors these keyphrases within the environment, thereby generating a set of grounding scores. These scores serve as the basis for the sequence-to-sequence model in CLIP-Nav, which computes the next set of actions based on the agent’s current state and grounding scores.
@@
To augment the efficacy of CLIP-Nav, the model incorporates a backtracking mechanism. This allows the agent to retrace its steps, facilitating revisions to prior decisions. Such a feature is particularly beneficial in the context of coarse-grained instruction-following tasks, wherein corrective backtracking may be essential. Additionally, the paper introduces a new performance metric, termed Relative Change in Success (RCS), to quantitatively assess generalizability in vision and language navigation tasks.
@@
c) SQA3D [80], as a benchmark, aims to assess the scene comprehension ability of embodied agents.The task necessitates that the agent garner an exhaustive understanding of its orientation within a 3D environment, guided by a text-based description, and subsequently generate precise answers to questions pertaining to that understanding.